Help me translate this into a quadratic equation and solve. Thanks!
Three consecutive even integers are such that the square of the third is 76 more than the square of the second. Find the three integers.

Answer by checkley77(12844)

LET X, X+2, & X+4 BE THE 3 CONSECUTIVE EVEN INTEGERS.
(X+4)^2=(X+2)^2+76
X^2+8X+16=X^2+4X+4+76
8X-4X=4+76-16=0
4X=64
X=64/4
X=16 ANSWER.
16+2=18 FOR THE MIDDLE NUMBER.
16+4=20 FOR THE THIRD NUMBER.
PROOF:
20^2=18^2+76
400=324+76
400=400